Determine the magnitude and direction of the electric force on the electron of a hydrogen atom exerted by the single proton that is the atom’s nucleus. Assume the average distance between the revolving electron and a proton is r= 0.53×10^-10m.
Answer #1

r = 0.53 * 10-10 m

a) F = k q1 q2 / r2

= (9 * 109 * (1.6 * 10-19)2) / (0.53 * 10-10)2

magnitude of the electric force = 8.202 * 10-8 N

b) direction is towards each other